OpenStudy (anonymous):

I NEED HELP PLEASE! simplify the following expression. Write the answer using scientific notation. (9x10^5) (3x10^10)

OpenStudy (calculusxy):

Multiply the constants (9 and 3) separately than the powers (10^5 and 10^10) \[(9 \times 3) \times (10^5 \times 10^{10})\]

OpenStudy (calculusxy):

@brunotbb

OpenStudy (anonymous):

Um would that be 1.2*10^16???

OpenStudy (anonymous):

@calculusxy

OpenStudy (calculusxy):

That's incorrect. First multiply \(9 \times 3\). What do you get?

OpenStudy (anonymous):

I don't understand these types of questions and 27

OpenStudy (anonymous):

OH is it 2.7*10^16???

OpenStudy (calculusxy):

Yes!

OpenStudy (anonymous):

Thank you so much ^_^ @calculusxy

OpenStudy (calculusxy):

You're welcome